### v2

# OmniGenome: Aligning RNA Sequences with Secondary Structures in Genomic Foundation Models

Abstract: The alignment between RNA sequences and structures in foundation models (FMs) has yet to be thoroughly investigated. Existing FMs have struggled to establish sequence-structure alignment, hindering the free flow of genomic information between RNA sequences and structures. In this study, we introduce OmniGenome, an RNA FM trained to align RNA sequences with respect to secondary structures based on structure-contextualised modelling. The alignment enables free and bidirectional mappings between sequences and structures by utilising the flexible RNA modelling paradigm that supports versatile input and output modalities, i.e., sequence and/or structure as input/output. We implement RNA design and zero-shot secondary structure prediction as case studies to evaluate the Seq2Str and Str2Seq mapping capacity of OmniGenome. Results on the EternaV2 benchmark show that OmniGenome solved of puzzles, whereas existing FMs only solved up to of the puzzles due to the oversight of sequence-structure alignment. We leverage four comprehensive in-silico genome modelling benchmarks to evaluate performance across a diverse set of genome downstream tasks, where the results show that OmniGenome achieves state-of-the-art performance on RNA and DNA benchmarks, even without any training on DNA genomes.
Keywords: RNA Foundation Model, RNA Design, Secondary Structure Prediction
1 Introduction
RNA is a critical type of molecule that encodes a vast array of biological regulatory elements that orchestrate crucial aspects of plant growth, development, and adaptation to environmental stresses. To decipher the genomic code in RNA and manipulate RNA engineering and design, current research mainly uses bioinformatics in solving RNA genome-oriented challenges. Recent advancements in large-scale pre-trained foundation models (FMs) have demonstrated their unprecedented potential to back up existing genome analysis, as FMs are capable of learning and predicting the complex ‘genomic language’ [1] hidden in genome encoding processes. Existing FMs have been widely employed as basic sequence feature extractors to improve the performance of diverse genome analysis tasks, such as secondary structure prediction [2, 3, 4, 5], degradation rate prediction [6, 7], and mRNA vaccine design [8, 9]. In RNA, it is intriguing that the functionality and stability are intertwined with its complex structures in molecular biology [10]. However, the role of the structure as a second ‘genomic language’ to interact with sequences and solve various RNA downstream tasks has been largely ignored.
Sequence-Structure Alignment in FMs
We define alignment between sequences and secondary structures111In this paper, all structures are referred to as the secondary structures. as the bidirectional information flows. Current FMs have been struggling to establish an alignment between RNA nucleotide sequences and their folded structures, thus impeding bidirectional genomic information flows. There has been a deep scientific challenge to align RNA sequences with structures because it is not deterministic to predict sequences from structures and vice versa. In other words, an identical sequence may be folded into different sub-optimal structures because the folding patterns of RNA sequences depend on various in-vivo factors [11]. Further, a structure can be folded from different sequences composed of variational combinations of nucleotide bases, as the example shown in Fig. 1. The oversight of such alignment in existing FMs causes outstanding issues in understanding and leveraging RNA structures, such as mRNA design. For example, recent state-of-the-art RNA FMs, RNA-FM [12] and RNA-MSM [13], only solved out of puzzles in in-silico RNA design [14]. This is because they fail to decipher corresponding sequences based on structures to guide RNA design.
To address the above two problems, we propose sequence-structure alignment in RNA FMs, which leverages the large-scale annotations of sequences and structures to build reliable structure to sequence (Str2Seq) and sequence to structure (Seq2Str) mappings, leading to an aligned FM dubbed OmniGenome. The sequence-structure alignment enables genomic information to freely flow between sequences and structures by introducing a flexible RNA modelling paradigm that supports versatile inputs and outputs modalities, i.e., sequence and/or structure as input/output. The sequence-structure alignment enables genomics information to freely flow between sequences and structures by introducing a flexible RNA modelling paradigm that supports versatile inputs and outputs modalities, i.e., sequence and/or structure as input/output. Furthermore, the sequence-structure alignment is designed to be architecture-agnostic and genome-agnostic. That is to say, it can be easily transferred to large-scale models with new architecture and different genome types like DNA.
Str2Seq Mapping
RNA structure serves as a vital input in most of the RNA genome analysis tasks. To induce the ability of Str2Seq mapping in genomic FMs, we formulate a structure-contextualised RNA sequence reconstruction task, which stems from the representation of RNA secondary structures in texts composed of dots and brackets. As the diagram is shown in Fig. 2, we first concatenate sequence-structure pairs as inputs and then mask a small portion of nucleotide bases in the sequence. Then, we pre-train OmniGenome to reconstruct the masked nucleotide bases given the structure contexts. This simple but effective formulation of Str2Seq mapping realises structure input awareness in genomics FM pre-training and provides substantial compatibility for structure-contextualised tasks, which has been verified in our RNA design benchmark.
Seq2Str Mapping
On the other hand, Seq2Str mapping, such as end-to-end secondary structure prediction (SSP) [15, 16], is another critical aspect of achieving the alignment. We generalise end-to-end structure pre-training [17] to OmniGenome pre-training. This large-scale structure pre-training on diversified genomes supervises OmniGenome to perform Seq2Str mapping. The problem of structure pre-training lies in RNA structure annotation scarcity, which leads to biased structure predictions [18] and barriers the structure prediction robustness on small datasets. To conduct Seq2Str mapping, tremendous secondary structure annotations are required to avoid data bias. A feasible solution to RNA structure pre-training is leveraging the plausible structures calculated based on the minimum free energy. In this paper, we leverage the popular ViennaRNA [19] to serve our purpose, ‘computing’ the structures for millions of RNA sequences and perform structure pre-training in OmniGenome.
Evaluation and Results
To validate the effectiveness of OmniGenome, we designed four large-scale genome benchmarks with diverse genomics tasks. The first one is the RNA genomics benchmark (RGB) compiled in the study, which contains diverse challenging genomics understanding tasks that benefit from the sequence-structure alignment, such as degradation rate prediction. The second benchmark is the plant genomics benchmark (PGB) [20] which contains millions of DNA sequences to evaluate the DNA sequence understanding tasks. In particular, we want to use this benchmark to evaluate the generalisability of OmniGenome among diversified species and genomes. The overall performance of OmniGenome (up to M parameters) on both two benchmarks consistently outperforms existing genomics FMs with up to improvement, even compared with Agro-NT [20] which contains billion parameters. The last two benchmarks, available in the appendix, are the genomics benchmark (GB) [21] and genomics understanding evaluation (GUE) [22], which serve as two additional DNA benchmarks to evaluate generalisability on non-plant genome modelling.
In addition, we also conduct zero-shot Seq2Str and Str2Seq prediction experiments to verify the performance of sequence-structure alignment. As revealed in the experiments in Sections 3.3 and 3.3, OmniGenome achieves up to a macro-F1 score in zero-shot Seq2Str prediction, i.e., secondary structure prediction, outperforming fine-tuned FMs and bioinformatics methods like ViennaRNA. In terms of Str2Seq prediction performance, we evaluate the performance of OmniGenome in the in-silico RNA design task. We solved of complex puzzles of the EternaV2 benchmark [14], while state-of-the-art FMs such as RNA-MSM and RNA-FM only solved up to . Besides, OmniGenome only takes less than one hour to solve most of the puzzles, while most RNA design methods need to take up to hours to solve even a single puzzle.
Open-source Toolkit and Tutorials
Open science is always the golden standard to promote this rising area of FM for genome modelling, which unfortunately lacks relevant high-quality resources such as code integrity, data availability, and pre-training pipeline. To address this gap, following the FAIR principles [23], we developed an open-source package222https://tinyurl.com/OmniGenome that includes step-by-step tutorials for FM pre-training and downstream tasks fine-tuning, to name a few. It provides ready-to-use genomics benchmarks and uses the API with only a few lines of code to streamline benchmarking purposes. We believe this will be a valuable resource to make this emerging AI for the RNA community to thrive.
2 Methodology
This section delineates the implementation details of OmniGenome including its entire pre-training workflow and downstream benchmarks.
2.1 RNA Tokenization for Alignment
We aim to implement a fine-grained alignment between RNA sequences and structures, where each base in the sequences reflects a structural label in ‘(’, ‘)’, ‘.’. Therefore, we propose an adapted implementation of the single nucleotide tokenization (SNT) method [1, 24] in OmniGenome, where the whole vocabulary, ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’, ‘(’, ‘)’, ‘.’, contains the nucleotide-level structural labels. We illustrate our tokenization based on an example shown in Fig. 3.
Our adapted SNT features bidirectional mappings between single nucleotide (SN) bases and structural labels required by sequence-structure alignment. Another reason for the adaption of SNT is that, in the realm of RNA genome modelling, the FM performance highly depends on the tokenization resolution [1, 24]. For example, the k-mers [25, 26] and BPE [27, 22] tokenization methods combine multiple bases into tokens and embeddings, which compromise modelling resolution and thus fail to the solution of fine-grained genomic tasks like structure prediction as well as base-level degrade rate prediction. Like other encoder-only models, e..g, BERT [27], we incorporated special tokens, e.g., ‘<mask>’, to implement masked language modelling.
2.2 Pre-training Objectives
As discussed in Section 1, a key desideratum for SN-level genome modelling is to build the alignment between RNA sequences with corresponding secondary structures. Bearing this in mind, we formulate two pre-training objectives, i.e., and , for Str2Seq and Seq2Str predictions, respectively. Besides, we aggregate these two objectives with the masked RNA language modelling objective MRLM to pre-train OmniGenome as follows:
where is the regularisation weight and represents the parameters of OmniGenome. The following paragraphs explain the design principles of each objective function used in equation (1).
is designed to enable OmniGenome to predict bases given structure-contextualised sequences with partially masked bases. This objective aims at Str2Seq tasks and teaches OmniGenome to interpret structure information and infer the masked sequences. To achieve this objective, we mask of the bases and structure tokens, encouraging the model to infer masked bases (i.e., ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’) and structure tokens (i.e., ‘(’, ‘)’, ‘.’). Specifically, is defined as the classic cross-entropy loss widely used in the masked language modelling:
(2) where is the number of masked nucleotide and structure tokens, and indicates the probability of predicting the masked nucleotide based on its context.
In terms of structure-out modelling, we implement to enable OmniGenome for Seq2Str predictions. Instead of directly feeding the secondary structure into OmniGenome as inputs, this objective employs the RNA secondary structures as labels for supervised training. This objective is implemented as a token-level classification, where the loss is defined in the following cross-entropy loss:
(3) where denotes the label of secondary structure at the -th position, and is the probability predicted by a linear classifier deployed on OmniGenome. is the length of an RNA sequence and denotes the number of the possible labels of structure, i.e., ‘(’, ‘)’, ‘.’.
The last objective, , is adapted to the conventional masked language modelling loss in NLP. It aims to improve the model’s understanding of genomic language in RNA sequences by predicting the masked or replaced of nucleotide bases. The definition of is similar to that of which only considers the prediction of masked bases rather than randomly replaced bases. The loss function of MRLM is well-known so we omit its formula here.
We cannot trust structure predictions (in ) while the structures are leaked in inputs (in ), i.e., the sequence inputs and outputs of these two objectives are exclusive. In practice, we only consider objectives either or for each input sequence. In the pre-training, of RNA sequences are used for the first two objectives, while the remaining are used for the latter two objectives. This proportion setting is concluded from our empirical experience to balance the capability of Str2Seq and Seq2Str predictions.
2.3 Model Architecture
OmniGenome adopts the Transformer encoder architecture with bidirectional multi-head attention. We do not adopt recent architectures like Mamba [28, 29] and Hyena [1] because our experiments in Table 5 and Table 6 show that these architectures are not competent at RNA genome understanding. This low performance is probably because RNA sequences are much shorter than DNA sequences in the wild.
We designed two variants, dubbed OmniGenome and OmniGenome with and million parameters respectively. Some key model specifications are summarised in Table 1.
| OmniGenome | 52M | 186M |
| # of Layers | ||
| Embedding dimension | ||
| Intermediate dimension | ||
| # of heads | ||
| # of parameters | M | M |
| Modelling length | ||
To improve the reproducibility of OmniGenome, we list the pre-training settings and hyperparameters as follows.
The learning rate is set to and the weight decay is set to .
We use AdamW as the optimiser with hyperparameters and .
We use a linear decay strategy with a warm-up period of steps in the learning rate scheduler.
The batch size is set to .
No dropout is applied during pre-training, and we use the rotary position embeddings [30] to further enhance the model’s scalability to long RNA sequences.
We built a distributed training environment with Nvidia RTX GPUs, while its configuration is introduced in Appendix 6. The pre-training was finished in approximately 1 and 3 weeks for OmniGenome and OmniGenome, respectively.
2.4 Pre-training Database: OneKP
Recent studies [24, 22] have shown that data diversity can enhance FM performance without significantly increasing model capacity. For the OmniGenome pre-training, we collected transcriptome data from the OneKP initiative333https://sites.google.com/a/ualberta.ca/onekp/ [31], which compiles large-scale RNA raw sequence database from plant species. The raw sequences are not available for pre-training before processing and filtering.
We adopt the following raw sequence data curation protocol to fit pre-training.
To enhance training efficiency and reduce bias, we removed all duplicate sequences.
To tackle incomplete transcriptome data and other noises, we discard sequences shorter than bases.
To facilitate the sequence-structure alignment training, we adopt ViennaRNA444https://github.com/ViennaRNA/ViennaRNA to obtain the secondary structures for the sequences.
2.5 Benchmark Suites
OmniGenome is designed as a general-purpose RNA FM that can be fine-tuned for a diverse set of downstream genomics predictive tasks. In this paper, we constructed a large-scale benchmark suite for RNA FMs. According to the category of genomes, we split the benchmark into two parts.
RNA Genomic Benchmark (RGB)
RGB is a collection of genome understanding tasks, as shown in Table 7. RGB contains SN-level tasks that are curated in this work or collected from published articles. The purpose of RGB is to benchmark FMs in challenging SN-level modelling tasks like the predictions of mRNA degradation rates and secondary structures. The sequence length in RGB ranges from to , which is enough for most RNA understanding tasks. These multi-species and SN-level tasks in RGB serve as the first comprehensive RNA benchmark to assess the modelling capabilities of FMs. For detailed information on each dataset, such as their sources and sizes, please refer to Appendix 9.1.
Plant Genomic Benchmark (PGB)
PGB555https://huggingface.co/datasets/InstaDeepAI/plant-genomic-benchmark [20] ) shown in Table 9 provides a large-scale and comprehensive suite of DNA genome tasks designed to evaluate the modelling capabilities of FMs in plant biology. PGB involves types of DNA downstream subtasks, including a range of critical tasks such as promoter strength prediction and gene expression regression. There are datasets in total with millions of DNA sequences to be evaluated in PGB, and the sequence lengths are up to , which is very long for most of the genomic FMs. Since the original evaluation protocol is not publicly available, we have re-implemented the auto-benchmark for all the subtasks from PGB in our package. By integrating diverse downstream tasks, PGB aims to facilitate the development of plant genomics and robust assessment. Due to computational limitations, we randomly sample a maximum of k examples in all downstream datasets in PGB to evaluate the FM’s performance.
2.6 Str2Seq Modelling Case: RNA Design
One of the most challenging practices addressed by OmniGenome is RNA design, which has not been settled in existing FMs because of the oversight of sequence-structure alignment. In this section, we provide a case of RNA design that requires the exploitation of sequence-structure alignment. To address RNA design, we introduce a simple but effective genetic algorithm (GA)666The experiment script is available at: https://tinyurl.com/OmniGenome/tutorials/zeroshot_omnigenome_RNA_design.py to search feasible RNA candidates, which is based on the Str2Seq prediction capability of OmniGenome. The GA implementation details main steps in the genetic algorithm and workflow visualisation are available in Appendix 10 and Fig. 5, respectively. In Section 3.3, the experimental results on the Eterna V2 dataset indicate an impressive performance in RNA design compared to existing methods.
3 Experiments
To evaluate the performance of OmniGenome across genome modelling, we implement experiments on diverse downstream tasks. We first evaluate the sequence-structure alignment capability of OmniGenome. Subsequently, we evaluate the overall performance of OmniGenome on two comprehensive genomic modelling benchmarks, i.e., RGB and the PGB, respectively. Finally, we include the GB and GUE in the appendix to evaluate the performance on non-plant genomes.
3.1 RNA Sequence Filtering
The pertaining involves RNA sequences and structures prediction, we take the data and annotation leakage problem seriously.
To avoid structure annotation leakage of downstream benchmarks, the secondary structure predictors for all FMs were randomly initialised for fair comparisons, which means the pre-trained structure predictor of OmniGenome was not used in benchmarks, except for zero-shot SSP experiments. Please find the source codes for details.
To reduce sequence leakage caused by evolutionary conservative sequences across multiple species, we use the ch-hit-est tool to calculate the sequence similarity between sequences from the OneKP database and downstream tasks. We adopt the similarity threshold of for ch-hit-est to eliminate sequences whose homogeneous sequences appeared in the OneKP database. Subsequently, we exploit the blastn tool to query potentially leaked sequences in downstream benchmark datasets and further alleviate the data leakage problem. The e-value has been set to for rigorous sequence filtering.
| Model | Tokenization | # of Params | Pre-training Data Scale | Pre-training Data Source | Species | Sequence Type |
| DNABERT-2 | BPE | M | B Tokens | The 1000 Genomes Project | Human + Species | DNA |
| NT-V2-M | k-mers | M | B Tokens | The 1000 Genomes Project, etc. | Human + Species | DNA |
| HyenaDNA-Large | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
| Caduceus | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
| Agro-NT-B | k-mers | M | B Tokens | Ensembl Plants Database | Edible Plants | DNA |
| SpliceBERT | SNT | M | M Sequences | UCSC Genome Browser | Multi-Vertebrates | precursor-mRNA |
| RNA-BERT | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
| RNA-MSM | SNT | M | RNA Families | The RNA Families Database | Multi-Species | ncRNA |
| RNA-FM | SNT | M | M Sequences | RNAcentral Database | Multi-Species | ncRNA |
| 3UTRBERT | k-mers | M | Sequences | The GENCODE Project | Human | mRNA ’UTR |
| OmniGenome | SNT | M | B Tokens | The OneKP Initiative | Plant Species | mRNA, CDS, UTR |
| OmniGenome | M |
3.2 Comparison Baselines
Apart from OmniGenome, we implement a plus variant, i.e., OmniGenome. In the context of OmniGenome, we assume the structure annotation from ViennaRNA is always available for enhancing the model based on structure-contextualised modelling. In SSP tasks, we can also use the ViennaRNA’s structure annotations as contexts to improve downstream SSP performance. Please refer to Appendix 8 for brief introductions of these FMs.
We can compare OmniGenome with the following RNA and DNA FMs shown in Table 2 as baselines to help evaluate the performance of OmniGenome. We are aware that some FMs are also developed for RNA, such as Uni-RNA [34], 5UTR-LM [35], etc. However, we cannot compare OmniGenome with them because their source codes are very hard to work with in our efforts or not publicly available.
3.3 Sequence-Structure Alignment Evaluation
In this section, we verify the sequence-structure alignment capability based on two experiments, i.e., Str2Seq prediction and zero-shot Seq2Str prediction via SSP and RNA design tasks, respectively. Overall, the results in Table 3 and Table 4 provide reliable evaluations of the FMs’ capabilities in sequence-structure alignment. This underscores OmniGenome’s efficacy in enabling genomic information to freely flow between structures and sequences.
RNA Design (Str2Seq) Evaluation
we demonstrate the Str2Seq prediction capability of OmniGenome based on RNA design. We employed the Eterna777https://github.com/eternagame/eterna100-benchmarking/blob/master/eterna100_vienna2.txt [14] V2 benchmark, which consists of specified secondary structures. This task aims to design RNA sequences based on reference structures. We develop a genetic algorithm (GA) which exploits masked nucleotide modelling (a.k.a., masked language modelling) to find plausible RNA sequences that solve RNA design puzzles. The implementation details can be found in Fig. 5 in Appendix 10. In the GA, the population size is set at , with iterations, and the mutation rate for each base is . The evaluation metric is accuracy following existing works which indicates the number of puzzles solved by FMs. The experimental results are available in Table 3.
| Model | Token. | EternaV2 (Acc) |
| RNAInverse | — | |
| 3UTRBERT | k-mers | |
| DNABERT2 | BPE | |
| SpliceBERT | SNT | |
| RNA-MSM | SNT | |
| RNA-FM | SNT | |
| OmniGenome | SNT | |
| OmniGenome | SNT |
We include a popular bioinformatics baseline of RNAInverse and select several recent DNA and RNA FMs which support masked language modelling. We exclude HyenaDNA in this experiment because it does not support masked nucleotide prediction. It is observed from Table 3 that RNAInverse solved of the RNA design puzzles, indicating a promising capability in RNA design. The FMs, such as 3UTRBERT and DNABERT2 fail in RNA design because they cannot handle SN-level modelling. Meanwhile, RNA-MSM, RNA-FM and SpliceBERT demonstrated trivial proficiency in RNA design, solving to puzzles. This observation suggests these FMs cannot precisely predict the bases without any Str2Seq prediction ability. With the help of Str2Seq, i.e., structure-contextualised sequence reconstruction, OmniGenome and OmniGenome significantly outperformed other FMs with and puzzles solved, respectively, underscoring the significance of Str2Seq in sequence-structure alignment. Besides, we expect an increase in performance with sufficient computational budgets and the findings provide crucial evidence of the significance of Str2Seq for RNA sequence design.
Zero-shot SSP (Seq2Str) Evaluation
This subsection evaluates both Seq2Str and Str2Seq prediction in sequence-structure alignment. The evaluation of Seq2Str is based on zero-shot SSP. We use OmniGenome and OmniGenome without fine-tuning to predict the secondary structures of sequences from the testing datasets and measure the macro-F1 score, where better structure prediction performance indicates a stronger capability for Seq2Str prediction. The experimental results are available in Table 4.
| Model | Zero-shot SSP (F1) | ||
| Archive2 | bpRNA | Stralign | |
| ViennaRNA | |||
| OmniGenome | |||
| OmniGenome | |||
| OmniGenome | |||
| OmniGenome | |||
| Model | PolyA | LncRNA | Chrom Acc | Prom Str | Term Str | Splice | Gene Exp | Enhancer |
| F1 | F1 | F1 | RMSE | RMSE | F1 | RMSE | F1 | |
| DNABERT2 | ||||||||
| HyenaDNA | ||||||||
| Caduceus | ||||||||
| NT-V2 | ||||||||
| Agro-NT | ||||||||
| SpliceBERT | ||||||||
| 3UTRBERT | ||||||||
| RNA-BERT | ||||||||
| RNA-MSM | ||||||||
| RNA-FM | ||||||||
| OmniGenome | ||||||||
| OmniGenome | ||||||||
| OmniGenome | ||||||||
| OmniGenome |
The results in Table 4 indicate that OmniGenome FMs mirrored the zero-shot secondary structure prediction (i.e., Seq2Str) performance of ViennaRNA. Moreover, OmniGenome and OmniGenome outperform OmniGenome FMs based on structure contexts from ViennaRNA. Given the ablation of structure contexts, OmniGenome also achieves performance comparable with ViennaRNA on the Archive2, bpRNA and RNAStralign datasets. Besides, we found that OmniGenome generally obtains better performance on a wide genome downstream tasks owing to the structure awareness, and random or noise structure contexts have no obvious effects on the structure prediction. We cannot compare with other FMs in the zero-shot SSP experiments, because existing FMs were not pertained for secondary structure prediction.
3.4 Results on RGB
| Model | mRNA | SNMD | SNMR | Archive2 | Stralign | bpRNA |
| RMSE | AUC | F1 | F1 | F1 | F1 | |
| ViennaRNA | N.A. | N.A. | N.A. | |||
| MXFold2 | N.A. | N.A. | N.A. | |||
| Ufold | N.A. | N.A. | N.A. | |||
| DNABERT2 | ||||||
| HyenaDNA | ||||||
| Caduceus | ||||||
| NT-V2 | ||||||
| Agro-NT | ||||||
| SpliceBERT | ||||||
| 3UTRBERT | ||||||
| RNABERT | ||||||
| RNA-MSM | ||||||
| RNA-FM | ||||||
| OmniGenome | ||||||
| OmniGenome | ||||||
| OmniGenome | ||||||
| OmniGenome |
The results in Table 6 demonstrate the performance of OmniGenome and its generalizability across various fine-grained RNA downstream tasks. It is observed that OmniGenome models achieve better results than both RNA and DNA FM baselines, including Agro-NT and DNABERT2, which contain hundreds of millions of parameters. This is because the existing FMs usually adopt k-mers or BPE tokenization that cannot handle SN resolution tasks, e.g., single nucleotide mutation detection and repair, and structure prediction. Because of the Seq2Str pre-training, OmniGenome and OmniGenome models exhibit strong results in secondary structure prediction, underscoring OmniGenome’s capabilities in SN-level RNA sequence understanding and manipulation.
3.5 Results on PGB
The PGB is a plant-oriented genomic benchmark. Although the benchmark datasets in PGB are DNA-based tasks, we can still evaluate the performance of OmniGenome and its generalizability on multi-modal (i.e., DNA and RNA) genomic tasks. The results in Table 5 reveal substantial variability in the performance of different FMs, where OmniGenome outperformed other baseline models across most tasks, particularly in tasks like Polyadenylation, Splice Site, and Enhancer Region classification, where they achieved the highest F1 scores. This suggests that OmniGenome’s architecture is particularly adept at handling complex genomic sequences. In comparison, existing FMs, e.g., NT-V2 and Agro-NT, showed lower performance with more parameters than OmniGenome. Besides, the performance of OmniGenome suggests that the structure context can further enhance the performance of genomic modelling. Overall, OmniGenome models achieve state-of-the-art performance on both benchmarks, especially for OmniGenome variants. The results underscore the importance of sequence-structure alignment in achieving complex genomic modelling tasks.
4 Related Works
Biological sequence modelling, including DNA, RNA, and protein, has attracted attention in recent years. Protein modelling, e.g., AlphaFold [36, 37, 38] and ESM [39], has been studied for many years compared to DNA and RNA modelling. In the realm of genomic sequence modelling, several early works aimed at addressing diversified genome downstream subtasks. For instance, DNABERT [40] adapts the architecture of BERT [27] for genomic sequence modelling, showing preliminary performance for in-silico genomic tasks. DNABERT2 [22], a multi-species FM improved based on DNABERT, proposes replacing k-mers tokenization with BPE tokenization to improve model performance. To explore the performance of large-scale FMs, the nucleotide transformers (V1 & V2) [26], AgroNT [20] and SegmentNT [41] leveraged billions of parameters to boost genomic sequence modelling and achieved promising performance in understanding DNA genome, with model scales up to billion and billion parameters, respectively. Agro-NT [20] was pre-trained on multi-species edible plant DNA sequences but failed to transfer effectively to RNA sequence modelling in our experiments. To address the modelling capacity problem caused by the remarkable lengths of genomes, there is a growing focus on the necessity of long-range sequence modelling and the introduction of autoregressive FMs, namely, HyenaDNA [1] and Evo [42]. In terms of RNA genomic modelling, there have been some preliminary works, such as scBERT [43], RNABERT [44], RNA-FM [12], RNA-MSM [45], and RNAErnie [46], to name a few. However, these methods have only trained the FMs on a limited-scale database, as RNA sequences are generally expensive to obtain. Some FMs focus on specific types of RNA sequences, such as coding sequences (CDS) [47], 5’ untranslated regions (5’UTR) [35], 3’ untranslated regions (3’UTR) [25], or precursor mRNA sequences [24], thus limiting the models’ ability to capture the diversity of RNA sequences. Uni-RNA [34] has been reported to achieve good performance due to the large scale of the model and database, however, it is not open-sourced and cannot be compared in the experiments.
In short, the existing RNA FMs neglect the significance of sequence-structure alignment in RNA genome modelling, while the 5UTR-LM [35] adopts the secondary structure prediction as a pre-training objective to achieve Seq2Str prediction in pre-training. However, these FMs are not available for Str2Seq mapping and suffer from limited model and data scales that fail to uncover the comprehensive efficacy of sequence-structure alignment on a wide set of genomic tasks. ERNIE-RNA [48] feeds the RNA structure along with the sequence into the model and improves the downstream tasks. However, it also ignores the significance of Str2Seq prediction capability. In a nutshell, existing FMs fail to achieve sequence-structure alignment without exception.
5 Conclusion
We introduced OmniGenome to tackle the challenge of sequence-structure alignment in genome modelling, which bridges the gap between sequence and structural information and improves the reliability of genome analysis. Experimental results on four comprehensive in-silico RNA and DNA benchmarks demonstrate that OmniGenome outperforms existing FMs across diversified downstream tasks, e.g., up to F1 score for SSP and accuracy of RNA design. The superior performance highlights the potential of sequence-structure alignment in the field of genomics.
Acknowledgment
[…]